### v1

# Combinatorial metaplexes and centrality indices for identifying higher-order interactions

Abstract
Simplicial complexes provide a combinatorial framework for modelling higher-order interactions (HoIs), that is, interactions involving more than two units. In this representation, simplices encode interactions and inclusion relations determine the structural organisation of the system. In classical graph-theoretic and simplicial complex-based models, vertices are treated as structurally homogeneous entities, and their intrinsic properties do not influence the formation of higher-dimensional simplices. However, in many real-world systems, HoIs are influenced by discrete intrinsic properties of the participating units. To address this limitation, we introduce the combinatorial metaplex, consisting of two interacting components: an underlying simplicial complex serving as an admissibility structure that specifies boundary-compatible higher-order simplex candidates, and a concentration layer that assigns concentration to vertices, extends it to simplices, and induces a threshold-based rule governing the realisation and weighting of higher-dimensional simplices. We investigate how simplex concentration influences the emergence of HoIs and identify the true HoIs as those satisfying both combinatorial admissibility and concentration-based realisation. Using facet-mediated adjacency, we define degree and closeness centrality for non-facet simplices. We further introduce a one-parameter family of centralities that interpolates between purely structural connectivity and concentration-induced coupling. This provides a framework for analysing structurally enriched higher-order networks.
Keywords: Combinatorial metaplex, Simplicial complex, Higher-order networks, Facet-mediated adjacency, Weighted walks, Centrality indices.
MSC: 05E45; 55U10; 05C82.
1  Introduction
Systems characterised by multiple interacting components and collective phenomena, including nonlinearity, emergence, and adaptive behaviour, are central to complex systems theory [25, 35, 52, 17]. These systems are often modelled using graph-theoretic representations, leading to complex networks in which system units are represented as vertices and their interactions as edges [32, 23, 47, 10, 14]. Examples include protein–protein interaction networks, brain networks, and ecological networks [32].
However, many complex systems exhibit interactions involving more than two units, in addition to pairwise interactions, referred to as higher-order interactions (HoIs) [4, 11, 16, 12]. To represent HoIs explicitly, more general combinatorial structures such as hypergraphs and simplicial complexes are commonly employed [39, 3, 46]. A hypergraph consists of a set of vertices together with a collection of subsets of these vertices, called hyperedges. Unlike a graph, where each edge contains exactly two vertices, a hyperedge may contain any number of vertices. A simplicial complex is a collection of subsets of a vertex set that is closed under inclusion: whenever a set belongs to the collection, all of its subsets also belong to it. The elements of a simplicial complex are called simplices, and a simplex containing $k+1$ vertices is called a $k$-simplex. In this representation, HoIs are modelled as simplices (such as triangles, tetrahedra, and higher-dimensional analogues), with the closure property ensuring that all lower-dimensional faces of an interaction are also present [30].
In many existing studies, representations of systems—such as graphs, hypergraphs, and simplicial complexes—provide powerful tools for modelling the interaction structure of a system [16, 10, 11]. However, these formalisms primarily capture patterns of interaction and typically disregard intrinsic properties of the individual units forming the system. In practice, internal attributes of system components often influence how interactions are formed and sustained [31].
The metaplex framework, introduced by Estrada and collaborators, unifies internal structure and interaction structure within a single representation [22, 31]. In this setting, each vertex carries an internal structure that supports a continuous dynamical process, while interactions between vertices are modelled by a network. This enables the study of diffusion processes evolving both within components and across their connections.
In many real-world systems, however, intrinsic properties of units are discrete and quantifiable rather than continuous [7, 6, 13]. Vertices may carry measurable attributes such as the $h$-index or citation count of an author in a co-authorship network, the population or economic capacity of a city in a transportation network, or the abundance of a species in an ecological corridor network [45, 43, 49]. These heterogeneous quantities influence interaction patterns but are not generated by internal dynamical processes. Studies on node-weighted networks demonstrate that such intrinsic vertex weights affect structural measures including centrality and flow [45].
Many higher-order network models employ abstract simplicial complexes, in which higher-dimensional simplices arise purely from combinatorial inclusion [52]. In particular, in clique-based constructions every complete subgraph is interpreted as a HoI, regardless of vertex-level attributes [35]. Consequently, HoIs emerge solely from pairwise closure, without incorporating the intrinsic concentration of the participating units.
This motivates the introduction of the combinatorial metaplex. In this framework, each vertex is endowed with an intrinsic structure representing the internal organisation of the corresponding unit. A concentration is defined on this structure as a scalar quantity capturing its intrinsic capacity. These vertex-level concentrations constitute the only fundamental numerical quantities in the model.
Unlike multilayer network models [15, 29], which represent different types of pairwise relations across distinct but interconnected network layers, the combinatorial metaplex introduces heterogeneity directly at the vertex level by assigning intrinsic structure to individual units.
Within this vertex-driven setting, higher-dimensional simplices are not assigned intrinsic weights. Instead, their weights are induced from the concentrations of their constituent vertices through the simplicial incidence structure, implemented via the fractional weight distribution principle [9, 34]. Because a vertex may participate in multiple simplices, its concentration is distributed among incident simplices according to this incidence relation. Consequently, any numerical value associated with a simplex arises solely from aggregated vertex contributions, and no independent weight is assigned at the simplex level. This distinguishes the combinatorial metaplex from weighted simplicial complexes (see, e.g., [44, 19, 5]), where a non-negative real-valued function is assigned directly to simplices, thereby prescribing numerical strengths to higher-order interactions.
In combinatorial metaplex, the realisation of a higher-order simplex is governed by a structural–quantitative condition. A simplex is realised only if the aggregated concentration of its boundary simplices exceeds a prescribed threshold, which acts as a global reference parameter. Thus, HoIs arise from two simultaneous requirements: combinatorial admissibility and sufficient inherited concentration. This ensures that realised HoIs are supported by intrinsic vertex-level properties rather than appearing solely through combinatorial closure.
Finally, we introduce centrality measures on combinatorial metaplexes — including degree, closeness, and harmonic centrality — in order to analyse how vertex-level intrinsic properties influence higher-order structural organisation.
A simplicial complex of dimension $1$ (i.e., the $1$-skeleton of $\Delta$) is a graph, and various centrality indices for pairwise networks have been introduced in the literature [48, 47, 33, 40]. The most basic centrality, degree centrality, counts the number of incident edges (that is, edges containing a vertex) to a vertex. In an edge-weighted graph, the strength of a vertex is defined as the aggregate weight of its incident edges. In classical graph-theoretic formulations, centrality measures have primarily focused on the edge-based adjacency structure. The incorporation of vertex weights into centrality definitions has been explored in [1, 45, 2, 28]. Our work extends these constructions by defining centrality measures for fully weighted graphs (with both vertex and edge weights), thereby providing a coherent treatment of intrinsic and adjacency-based properties.
Several notions of adjacency between simplices have been proposed in the literature. Centrality in higher-order networks provides a quantitative framework for identifying the most influential simplices (or vertices). Extending graph-based centralities to simplicial complexes requires an appropriate notion of adjacency between simplices. The adjacency between two distinct $k$-simplices has been defined in terms of upper adjacency (when there exists a $(k+1)$-simplex containing both) and lower adjacency (when there exists a (k1)(k-1)-simplex contained in both) [26]. A generalised notion of adjacency, in which two distinct $k$-simplices are lower adjacent but not upper adjacent, was introduced in [24]. The adjacency between arbitrary simplices were defined in [41, 42].
The generalised degree centrality in $d$-exclusive simplicial complex of $\Delta$ (i.e.,family of $d$-simplices of $\Delta$ and their faces) was defined in [20] and subsequently extended to betweenness and closeness centralities in [36]. Using a tensorial representation of the $d$-exclusive simplicial complex, the strength of a simplex in a weighted simplicial complex was defined in [21]. In defining centralities for the combinatorial metaplex, we follow the facet-based adjacency framework proposed in [37], in which two distinct $k$-simplices are said to be adjacent whenever they are contained in a common facet. Characterising simplices with maximal centrality yields structural insight into the combinatorial metaplex and consequently, into the underlying system it models.
The main contributions of this work are as follows: (1) We introduce the combinatorial metaplex, a discrete two-layered construction that couples intrinsic vertex-level data with higher-order simplicial structure. (2) We extend the vertex concentration map to the entire simplicial complex via a fractional contribution map, propagating intrinsic vertex properties to higher-dimensional simplices through the incidence structure. (3) We establish a conservation relation between vertex-level concentrations and maximal higher-order simplices: the total induced weight of all facets coincides exactly with the total concentration assigned to the 0-simplices. (4) We develop a concentration-based framework for reconstructing simplicial complexes from data. Unlike clique-based constructions, which treat every clique of size greater than two as a HoI, our approach identifies only those simplices whose aggregated boundary concentration exceeds a prescribed threshold, thereby detecting concentration-supported HoIs. (5) We define degree centrality for fully weighted graphs consistent with previous definitions in weighted graphs. (6) We introduce degree, closeness, and harmonic centralities for the combinatorial metaplex using a facet-mediated notion of adjacency and weighted walks.
Section 2 presents the standard definitions from combinatorial algebraic topology required for completeness. These notions are well established in the literature [18, 27, 30, 36, 35].
2  Preliminaries
Let $V=\{v_{0},v_{1},\ldots,v_{n}\}$ be a finite, non-empty set of vertices, where $n\in\mathbb{N}$. We denote the power set of $V$ as $\mathcal{P}(V)$.
Definition 2.1.
A graph $G$ is a three-tuple $G=(V,E,\psi_{G})$, where:
  1. (i)
    $V$ is the set of vertices,
  2. (ii)
    $E$ is the set of edges,
  3. (iii)
    $\psi_{G}:E\rightarrow V\cup[V]^{2}$ is the incidence map that assigns to each edge either one or two vertices (its endpoints), where $V]^{2}=\{X\in\mathcal{P}(V)\mid|X|=2\}$.
Definition 2.2.
A simple graph is a graph $G=(V,E)$, where $E\subseteq[V]^{2}$. That is, each edge is a two-element subset of $V$; in particular, there are no loops or multiple edges.
Throughout this work, unless otherwise specified, the term graph refers exclusively to a simple graph.
Definition 2.3.
Let $G=(V,E)$ be a graph and let $Z\subseteq V$. The induced subgraph of $G$ on $Z$, denoted by $G[Z]$, is the graph whose vertex set is $Z$ and whose edge set is
$E_{Z}=\big\{\{u,v\}\in E\;\big|\;u,v\in Z\big\}.$
That is, $G[Z]$ contains exactly those edges of $G$ whose endpoints both lie in $Z$.
Definition 2.4.
A simplex over $V$ is a non-empty subset $\sigma\subseteq V$. If the cardinality $|\sigma|=d+1$, the dimension of $\sigma$ is defined as
$dim(\sigma)=|\sigma|-1=d.$
In this case, $\sigma$ is called a $d$-simplex, and is often denoted as $\sigma^{(d)}$.
Definition 2.5.
Let $\sigma^{(q)}$ be a $q$-simplex over a vertex set $V$. The boundary of $\sigma^{(q)}$, denoted $\mathcal{B}(\sigma^{(q)})$, is the family of all (q1)(q-1)-dimensional faces of $\sigma^{(q)}$.
Formally, we define
$\mathcal{B}(\sigma^{(q)})=\left\{\tau\subsetneq\sigma^{(q)}\mid dim(\tau)=q-1\right\}.$
Definition 2.6.
If $\sigma^{(p)}\subseteq\sigma^{(q)}$, then $\sigma^{(p)}$ is called a $p$-dimensional face of $\sigma^{(q)}$. If $p<q$, then $\sigma^{(p)}$ is called a proper face.
Definition 2.7.
A collection $\Delta$ of simplices is called an abstract simplicial complex if it satisfies the condition:
$\forall\sigma\in\Delta,\text{ and }(\tau\in\mathcal{P}(V),\text{ if }\tau\subseteq\sigma\Rightarrow\tau\in\Delta).$
That is, every face of a simplex in $\Delta$ must also belong to $\Delta$. The vertex set of $\Delta$ is denoted $V(\Delta)=\{v_{i}\in V\mid\{v_{i}\}\in\Delta\}$.
In this terminology, a 0-simplex is a vertex, a $1$-simplex is an edge, a $2$-simplex is a triangle, a $3$-simplex is a tetrahedron, and, in general, a $k$-simplex consists of $k+1$ vertices. Throughout this work, the term simplicial complex refers to an abstract simplicial complex.
Definition 2.8.
The dimension of a simplicial complex is the maximum number $q\in\mathbb{N}$ such that $\Delta$ contains a $q$-simplex:
$dim(\Delta)=\max\{q\in\mathbb{N}\mid\sigma^{(q)}\in\Delta\}.$
Definition 2.9.
The collection of all simplices in $\Delta$ of dimension at most $d$ is called the $d$-skeleton of $\Delta$, and is denoted by:
$\Delta^{(d)}=\{\sigma\in\Delta\mid dim(\sigma)\leq d\}.$
The 1-skeleton of a simplicial complex $\Delta$ is the graph formed by its 0- and 1-simplices. The vertices of the graph correspond to the 0-simplices, and an edge exists between two vertices if and only if the corresponding 1-simplex belongs to the 1-skeleton of $\Delta$.
Definition 2.10.
A facet of a simplicial complex $\Delta$ is a simplex that is not contained in any other simplex of $\Delta$. Formally, a simplex $\gamma\in\Delta$ is a facet if
$\nexists\tau\in\Delta\setminus\{\gamma\}\text{ such that }\gamma\subseteq\tau$
Definition 2.11.
The set of $q$-simplices of $\Delta$ is denoted by $S_{q}$:
$S_{q}=\{\sigma\in\Delta\mid dim(\sigma)=q\}\qquad\text{with}\qquad|S_{q}|=n_{q}.$
We denote an enumeration of all $q$-dimensional simplices of $\Delta$ as:
$S_{q}=\{\sigma_{1}^{(q)},\sigma_{2}^{(q)},\ldots,\sigma_{n_{q}}^{(q)}\}.$
Next, we define the upper degree following the framework in [26].
Definition 2.12.
Let $0\leq q<dim(\Delta)$. The upper degree of a $q$-simplex $\sigma_{i}^{(q)}$ is defined as the number of $(q+1)$-simplices containing it:
$k_{\sigma_{i}^{(q)}}=\left|\left\{\sigma^{(q+1)}\in S_{q+1}\mid\sigma_{i}^{(q)}\subsetneq\sigma^{(q+1)}\right\}\right|.$
Definition 2.13.
A graph $G=(V,E)$ is said to be complete if
$E(G)=\big\{\{u,v\}\subseteq V\;\big|\;u\neq v\big\}.$
Definition 2.14.
Let $G=(V,E)$ be a graph. The clique complex of $G$, denoted by $C(G)$, is the abstract simplicial complex defined by
$C(G)=\big\{\sigma\subseteq V\;\big|\;G[\sigma]\text{ is a complete graph}\big\}.$
Definition 2.15.
A metaplex is a $4$-tuple
$\mathcal{X}=(V,E,\mathcal{I},\omega),$
where $(V,E)$ is a graph,
$\omega=\{\Omega_{j}\}_{j=1}^{k}$
is a set of locally compact metric spaces $\Omega_{j}$ with Borel measures $\mu_{j}$, and
$\mathcal{I}:V\to\omega$
is an assignment map [22].
3  The combinatorial metaplex and the weight map
3.1  The combinatorial metaplex
As mentioned previously, the metaplex framework introduced in [22], models the internal vertex structure using locally compact metric–measure spaces. These spaces naturally support continuous dynamics, such as diffusion. However, in many systems, the internal structure of units is inherently discrete and does not admit a meaningful continuous description. In such cases, replacing internal metric–measure spaces by a finite combinatorial structure provides a natural generalisation.
When vertices possess discrete intrinsic properties that affect the global topology of a system, and these properties can be expressed as a finite discrete combinatorial structure, we introduce the discrete metaplex framework. Within this framework, each vertex is associated with an internal discrete structure endowed with a discrete measure, which we formalise by defining a concentration map on the set of vertices.
Definition 3.1.
A combinatorial metaplex (CM) is a quintuple
$\mathcal{K}_{c}=(V,\Delta,\mathcal{Y},\mathcal{W},\mathcal{I}),$ (1)
where:
  1. (i)
    $V=\{v_{0},v_{1},\dots,v_{n}\}$ is a finite, non-empty set of vertices.
  2. (ii)
    $\Delta$ is a finite simplicial complex over $V$.
  3. (iii)
    $\mathcal{Y}=\{Y_{i}\}_{i=0}^{n}$ is a family of finite, non-empty sets indexed by the vertices.
  4. (iv)
    $\mathcal{W}=\{w_{i}\}_{i=0}^{n}$ is a collection of non-negative weight maps
    $w_{i}:Y_{i}\to\mathbb{Q}\setminus\mathbb{Q}^{-},$
    such that $w_{i}$ is not identically zero, i.e.,
    $\sum_{x\in Y_{i}}w_{i}(x)>0\qquad\text{for all }0\leq i\leq n.$
  5. (v)
    $\mathcal{I}:V\to\mathcal{Y}$ is an assignment map defined by
    $\mathcal{I}(v_{i})=Y_{i}\qquad\text{for all }0\leq i\leq n.$
Each $w_{i}$ induces a discrete measure $\mu_{i}$ on the power set $\mathcal{P}(Y_{i})$, such that for any subset $A\subseteq Y_{i}$,
$\mu_{i}(A)=\sum_{x\in A}w_{i}(x).$
Thus, each triple $(Y_{i},\mathcal{P}(Y_{i}),\mu_{i})$ defines a discrete measure space associated uniquely with the vertex $v_{i}\in V$ [38].
The concentration at a vertex $v\in V$ is defined as the discrete measure of its internal structure:
Definition 3.2.
Given a CM as in (1), for a vertex $v_{j}\in V,\ 0\leq i\leq n$, let $\mathcal{I}(v_{j})=Y_{j}\in\mathcal{Y}$. The concentration is a map
$a:V\to\mathbb{Q}^{+},$
defined by
$a(v_{j}):=\mu_{j}(Y_{j})=\sum_{x\in Y_{j}}w_{j}(x).$
In the CM framework, vertices are endowed with a combinatorial structure that encodes their intrinsic discrete properties. This internal structure is represented by strictly positive rational values. By assigning positive weights, we ensure that the contribution of each vertex is strictly positive and is inherited by higher-dimensional simplices in the hierarchical extension.
Henceforth, the term ‘weight’ will also be used to denote concentration. In the subsequent sections, we develop a method to extend the weight map from the vertex set to the simplicial complex.
3.2  The weight map on 0-simplices
The set of vertices $V$ constitutes the primitive objects of the simplicial complex. In the classical theory of simplicial complexes, these vertices are treated as structurally indistinguishable. In the CM, however, this homogeneity is broken: vertices are differentiated through the assignment of positive rational values, representing their internal properties, via the concentration map.
Let $\Delta$ be a simplicial complex over a vertex set $V$. We define the weight map as follows:
$a:V\to\mathbb{Q}^{+},\hskip 17.00024ptv\mapsto a(v).$
Restricting the map $a$ to the active vertex set of the complex $\Delta$, we obtain the following:
$a|_{V(\Delta)}:V(\Delta)\to\mathbb{Q}^{+}.$
This restriction naturally induces a map on the set of 0-simplices (vertices) of $\Delta$. Denoting this map by $a^{\prime}$, we have:
$a^{\prime}:S_{0}\to\mathbb{Q}^{+},\hskip 17.00024pt\{v_{i}\}\mapsto a^{\prime}(\{v_{i}\})=a|_{V(\Delta)}(v_{i}).$
For convenience, we again denote the map $a^{\prime}$ by $a$, unless this causes ambiguity in notation. With the vertices now distinguished by their concentration values, we next define how this weight extends to higher-dimensional simplices (edges, triangles, etc.), thereby yielding a fully weighted simplicial structure.
3.3  The fractional weight contribution map and contribution numbers
In this section, we describe a constructive inductive procedure for extending the weight map dimension by dimension over the simplicial complex. The weights are initially assigned to the vertices, and the weight of a $q$-simplex is determined by the weights of its boundary simplices.
Let the weight map on the set of 0-simplices, $S_{0}$, be
$a:S_{0}\to\mathbb{Q}^{+}.$
We begin by defining the fractional weight contribution map $a_{1}$, which assigns to each pair of a 0-simplex and a $1$-simplex a non-negative rational number representing the fraction of the weight of the vertex that contributes to the edge.
Definition 3.3.
The fractional weight contribution map $a_{1}$ is defined as
$a_{1}:S_{0}\times S_{1}\longrightarrow\mathbb{Q}\setminus\mathbb{Q}^{-},$
$(\sigma^{(0)},\sigma^{(1)})\longmapsto a_{1}(\sigma^{(0)},\sigma^{(1)}),$
satisfying the following conditions:
  • (i)
    If $\sigma^{(0)}\not\subset\sigma^{(1)}$, then $a_{1}(\sigma^{(0)},\sigma^{(1)})=0.$
  • (ii)
    If $\sigma^{(0)}\subset\sigma^{(1)}$, then $0<a_{1}(\sigma^{(0)},\sigma^{(1)})\leq 1.$
  • (iii)
    For each $\sigma^{(0)}\in S_{0}$ with $k_{\sigma^{(0)}}\neq 0$, $\sum_{\sigma^{(1)}\in S_{1}}a_{1}(\sigma^{(0)},\sigma^{(1)})=1.$
From properties (i) and (ii), it follows that
$0\leq a_{1}(\sigma^{(0)},\sigma^{(1)})\leq 1\qquad\text{for all }\sigma^{(0)}\in S_{0},\ \sigma^{(1)}\in S_{1}.$
If $k_{\sigma^{(0)}}=0$, then by property (i),
$a_{1}(\sigma^{(0)},\sigma^{(1)})=0\qquad\text{for all }\sigma^{(1)}\in S_{1}.$
Based on this fractional map, we associate to each pair $(\sigma^{(0)},\sigma^{(1)})$ a contribution number $n^{\sigma^{(1)}}_{\sigma^{(0)}}$, defined by
$n^{\sigma^{(1)}}_{\sigma^{(0)}}=a_{1}(\sigma^{(0)},\sigma^{(1)})\,a(\sigma^{(0)}).$
This quantity represents the absolute weight contribution of the vertex $\sigma^{(0)}$ to the edge $\sigma^{(1)}$. Suppose that a 0-simplex $\sigma^{(0)}$ is contained in $k_{\sigma^{(0)}}\neq 0$ distinct $1$-simplices
$\sigma^{(1)}_{i_{1}},\ \sigma^{(1)}_{i_{2}},\ \ldots,\ \sigma^{(1)}_{i_{k_{\sigma^{(0)}}}},\qquad 1\leq i_{j}\leq n_{1},\ 1\leq j\leq k_{\sigma^{(0)}}.$
Then, the weight $a(\sigma^{(0)})$ is partitioned among these $k_{\sigma^{(0)}}$ simplices, and by property (ii),
$0<n^{\sigma^{(1)}_{i_{j}}}_{\sigma^{(0)}}\leq a(\sigma^{(0)}),\qquad 1\leq j\leq k_{\sigma^{(0)}}.$
Moreover, by property (iii), the conservation property holds:
$\sum_{j=1}^{k_{\sigma^{(0)}}}n^{\sigma^{(1)}_{i_{j}}}_{\sigma^{(0)}}=a(\sigma^{(0)}).$
For $1$-simplices not containing $\sigma^{(0)}$, or when $k_{\sigma^{(0)}}=0$, property (i) implies that
$n^{\sigma^{(1)}}_{\sigma^{(0)}}=0.$
Consequently,
$\sum_{\sigma^{(1)}\in S_{1}}n^{\sigma^{(1)}}_{\sigma^{(0)}}=a(\sigma^{(0)}),$
so the total weight of each vertex is conserved.
For convenience, we also denote the extended map on $S_{1}$ again by $a$, unless this causes ambiguity in notation. Finally, the weight of a $1$-simplex $\sigma^{(1)}\in S_{1}$ is defined as the sum of the contributions of its constituent vertices. Thus, extending the weight map $a$ to $S_{1}$, we obtain
$a(\sigma^{(1)})=\sum_{i=1}^{n_{0}}n_{\sigma^{(0)}_{i}}^{\sigma^{(1)}}=\sum_{i=1}^{n_{0}}a_{1}(\sigma^{(0)}_{i},\sigma^{(1)})\,a(\sigma^{(0)}_{i})$,
by property $(ii)$, $a(\sigma^{(1)})>0,\ \forall\sigma^{(1)}\in S_{1}$.
3.4  The general fractional weight contribution map and higher-order extensions
We generalise the previous construction to obtain an inductive procedure for extending the weight map from (q1)(q-1)-simplices to $q$-simplices.
To define the extension of the weight map from $S_{q-1}$ to $S_{q}$, we introduce the fractional weight contribution map $a_{q}$, which determines how the weight of a (q1)(q-1)-simplex is distributed among the $q$-simplices that contain it.
The value $a_{q}(\sigma^{(q-1)},\sigma^{(q)})$ represents the fraction of the weight of the (q1)(q-1)-simplex $\sigma^{(q-1)}$ contributed to the weight of the $q$-simplex $\sigma^{(q)}$.
Definition 3.4.
Let $1\leq q\leq dim(\Delta)$. The fractional weight contribution map $a_{q}$ is defined as
$a_{q}:S_{q-1}\times S_{q}\longrightarrow\mathbb{Q}\setminus\mathbb{Q}^{-},$
$(\sigma^{(q-1)},\sigma^{(q)})\longmapsto a_{q}(\sigma^{(q-1)},\sigma^{(q)}),$
satisfying the following conditions:
  • (i)
    If $\sigma^{(q-1)}\not\subset\sigma^{(q)}$, then $a_{q}(\sigma^{(q-1)},\sigma^{(q)})=0.$
  • (ii)
    If $\sigma^{(q-1)}\subset\sigma^{(q)}$, then $0<a_{q}(\sigma^{(q-1)},\sigma^{(q)})\leq 1.$
  • (iii)
    For each $\sigma^{(q-1)}\in S_{q-1}$ with $k_{\sigma^{(q-1)}}\neq 0$, $\sum_{\sigma^{(q)}\in S_{q}}a_{q}(\sigma^{(q-1)},\sigma^{(q)})=1.$
From (i) and (ii), we obtain
$0\leq a_{q}(\sigma^{(q-1)},\sigma^{(q)})\leq 1\qquad\text{for all }\sigma^{(q-1)}\in S_{q-1},\ \sigma^{(q)}\in S_{q}.$
If $k_{\sigma^{(q-1)}}=0$, then by property (i),
$a_{q}(\sigma^{(q-1)},\sigma^{(q)})=0\qquad\text{for all }\sigma^{(q)}\in S_{q}.$
Based on this map, we define a contribution number
$n^{\sigma^{(q)}}_{\sigma^{(q-1)}}=a_{q}(\sigma^{(q-1)},\sigma^{(q)})\,a(\sigma^{(q-1)}),$
which represents the absolute weight contribution of the (q1)(q-1)-simplex $\sigma^{(q-1)}$ to the $q$-simplex $\sigma^{(q)}$.
Suppose that a (q1)(q-1)-simplex $\sigma^{(q-1)}$ is contained in $k_{\sigma^{(q-1)}}\neq 0$ distinct $q$-simplices
$\sigma^{(q)}_{i_{1}},\ \sigma^{(q)}_{i_{2}},\ \ldots,\ \sigma^{(q)}_{i_{k_{\sigma^{(q-1)}}}}\qquad 1\leq i_{j}\leq n_{q},\ 1\leq j\leq k_{\sigma^{(q-1)}}.$
Then the weight $a(\sigma^{(q-1)})$ is partitioned among these simplices, and by property (ii),
$0<n^{\sigma^{(q)}_{i_{j}}}_{\sigma^{(q-1)}}\leq a(\sigma^{(q-1)}),\qquad 1\leq j\leq k_{\sigma^{(q-1)}},$
with the conservation property
$\sum_{j=1}^{k_{\sigma^{(q-1)}}}n^{\sigma^{(q)}_{i_{j}}}_{\sigma^{(q-1)}}=a(\sigma^{(q-1)}).$
For $q$-simplices not containing $\sigma^{(q-1)}$, or when $k_{\sigma^{(q-1)}}=0$, we get
$n^{\sigma^{(q)}}_{\sigma^{(q-1)}}=0.$
Consequently,
$\sum_{\sigma^{(q)}\in S_{q}}n^{\sigma^{(q)}}_{\sigma^{(q-1)}}=a(\sigma^{(q-1)}),$
so the total weight of each (q1)(q-1)-simplex is conserved.
Finally, the weight of a $q$-simplex $\sigma^{(q)}\in S_{q}$ is defined as the sum of the contributions of its boundary simplices:
$a(\sigma^{(q)})=\sum_{\sigma^{(q-1)}\in S_{q-1}}n^{\sigma^{(q)}}_{\sigma^{(q-1)}}=\sum_{\sigma^{(q-1)}\in S_{q-1}}a_{q}(\sigma^{(q-1)},\sigma^{(q)})\,a(\sigma^{(q-1)}),$
by property $(ii)$, $a(\sigma^{(q)})>0,\ \forall\sigma^{(q)}\in S_{q}$.
Thus, inductively, the concentration map $a$ extends from $S_{0}$ to all simplices of $\Delta$, yielding a well-defined weight map
$a:\Delta\longrightarrow\mathbb{Q}\setminus\mathbb{Q}^{-}.$
We now define the composition of two consecutive fractional weight contribution maps.
3.5  The composition of the fractional weight contribution maps
The composition of two consecutive fractional weight contribution maps is again a fractional weight contribution map, describing the fraction of the weight of a (q1)(q-1)-simplex contributed to a $(q+1)$-simplex. This composed map enables the computation of weight contributions across multiple dimensions, in particular from vertices to higher-dimensional simplices.
The composition map is useful for determining the fractional contribution number of a 0-simplex to any $q$-simplex.
To derive this composition map, we start with the definition of $a(\sigma^{(q+1)})$ and expand it iteratively:
$a(\sigma^{(q+1)})$ $=\sum_{i=1}^{n_{q}}a_{q+1}\bigl(\sigma_{i}^{(q)},\sigma^{(q+1)}\bigr)\,a(\sigma_{i}^{(q)})$
Substituting the value of $a(\sigma_{i}^{(q)})$ from the previous dimension:
$=\sum_{i=1}^{n_{q}}a_{q+1}\bigl(\sigma_{i}^{(q)},\sigma^{(q+1)}\bigr)\left(\sum_{j=1}^{n_{q-1}}a_{q}\bigl(\sigma_{j}^{(q-1)},\sigma_{i}^{(q)}\bigr)\,a(\sigma_{j}^{(q-1)})\right)$
$=\sum_{i=1}^{n_{q}}\sum_{j=1}^{n_{q-1}}a_{q+1}\bigl(\sigma_{i}^{(q)},\sigma^{(q+1)}\bigr)\,a_{q}\bigl(\sigma_{j}^{(q-1)},\sigma_{i}^{(q)}\bigr)\,a(\sigma_{j}^{(q-1)})$
Interchanging the order of summation (valid as all summand terms are positive):
$=\sum_{j=1}^{n_{q-1}}\left(\sum_{i=1}^{n_{q}}a_{q+1}\bigl(\sigma_{i}^{(q)},\sigma^{(q+1)}\bigr)\,a_{q}\bigl(\sigma_{j}^{(q-1)},\sigma_{i}^{(q)}\bigr)\right)a(\sigma_{j}^{(q-1)}).$
Thus we may define,
$a(\sigma^{(q+1)})=\sum_{j=1}^{n_{q-1}}(a_{q+1}\circ a_{q})\bigl(\sigma_{j}^{(q-1)},\sigma^{(q+1)}\bigr)\,a(\sigma_{j}^{(q-1)}),$
where
$(a_{q+1}\circ a_{q})\bigl(\sigma^{(q-1)},\sigma^{(q+1)}\bigr)=\sum_{\sigma^{(q)}\in S_{q}}a_{q+1}\bigl(\sigma^{(q)},\sigma^{(q+1)}\bigr)\,a_{q}\bigl(\sigma^{(q-1)},\sigma^{(q)}\bigr).$
Theorem 3.1.
Let $1\leq q\leq\dim(\Delta)-1$. Let $a_{q}$ and $a_{q+1}$ be the fractional weight contribution maps. Then their composition map is
$a_{q+1}\circ a_{q}:S_{q-1}\times S_{q+1}\longrightarrow\mathbb{Q}\setminus\mathbb{Q}^{-},$
$(\sigma^{(q-1)},\sigma^{(q+1)})\longmapsto(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)}),$
where
$(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})=\sum_{\sigma^{(q)}\in S_{q}}a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})a_{q}(\sigma^{(q-1)},\sigma^{(q)}).$
The map $a_{q+1}\circ a_{q}$ satisfies the following conditions:
  • (i)
    If $\sigma^{(q-1)}\not\subset\sigma^{(q+1)}$, then $(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})=0.$
  • (ii)
    If $\sigma^{(q-1)}\subset\sigma^{(q+1)}$, then $0<(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})\leq 1.$
  • (iii)
    For each $\sigma^{(q-1)}\in S_{q-1}$ if $\exists\sigma^{(q+1)}\in S_{q+1}$ such that $\sigma^{(q-1)}\subset\sigma^{(q+1)}$, then
    $\sum_{\sigma^{(q+1)}\in S_{q+1}}(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})=1.$
Proof.
Since
$a_{q}(\sigma^{(q-1)},\sigma^{(q)})\in\mathbb{Q}\setminus\mathbb{Q}^{-},\hskip 17.00024pta_{q+1}(\sigma^{(q)},\sigma^{(q+1)})\in\mathbb{Q}\setminus\mathbb{Q}^{-},$
their product belongs to $\mathbb{Q}\setminus\mathbb{Q}^{-}$, and hence
$(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})\in\mathbb{Q}\setminus\mathbb{Q}^{-}.$
Proof of (i): If $\sigma^{(q-1)}\not\subset\sigma^{(q+1)}$, then there does not exist $\sigma^{(q)}\in S_{q}$ such that
$\sigma^{(q-1)}\subset\sigma^{(q)}\subset\sigma^{(q+1)}.$
Hence for every $\sigma^{(q)}\in S_{q}$, either
$a_{q}(\sigma^{(q-1)},\sigma^{(q)})=0\qquad\text{or}\qquad a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})=0.$
Therefore every summand in
$(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})=\sum_{\sigma^{(q)}\in S_{q}}a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})a_{q}(\sigma^{(q-1)},\sigma^{(q)})$
is zero, and thus
$(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})=0.$
Proof of (ii): If $\sigma^{(q-1)}\subset\sigma^{(q+1)}$, then there exists at least one $\sigma^{(q)}\in S_{q}$ such that
$\sigma^{(q-1)}\subset\sigma^{(q)}\subset\sigma^{(q+1)}.$
For such $\sigma^{(q)}$,
$a_{q}(\sigma^{(q-1)},\sigma^{(q)})>0\qquad\text{and}\qquad a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})>0.$
Hence at least one summand in the defining sum is strictly positive, and therefore
$0<(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)}).$
Since
$0\leq a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})\leq 1\qquad\text{and}\qquad a_{q}(\sigma^{(q-1)},\sigma^{(q)})\geq 0,$
we have
$a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})\,a_{q}(\sigma^{(q-1)},\sigma^{(q)})\leq a_{q}(\sigma^{(q-1)},\sigma^{(q)}).$
Summing over $\sigma^{(q)}\in S_{q}$ gives
$(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})$ $=\sum_{\sigma^{(q)}\in S_{q}}a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})a_{q}(\sigma^{(q-1)},\sigma^{(q)})$
$\leq\sum_{\sigma^{(q)}\in S_{q}}a_{q}(\sigma^{(q-1)},\sigma^{(q)})$
$=1.$
Proof of (iii): Let $\sigma^{(q-1)}\in S_{q-1}$ such that there exists $\sigma^{(q+1)}\in S_{q+1}$ with $\sigma^{(q-1)}\subset\sigma^{(q+1)}$. Then
$\sum_{\sigma^{(q+1)}\in S_{q+1}}(a_{q+1}\circ a_{q})(\sigma^{(q-1)},\sigma^{(q+1)})$ $=\sum_{\sigma^{(q+1)}\in S_{q+1}}\sum_{\sigma^{(q)}\in S_{q}}a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})a_{q}(\sigma^{(q-1)},\sigma^{(q)})$
$=\sum_{\sigma^{(q)}\in S_{q}}\left(\sum_{\sigma^{(q+1)}\in S_{q+1}}a_{q+1}(\sigma^{(q)},\sigma^{(q+1)})\right)a_{q}(\sigma^{(q-1)},\sigma^{(q)})$
$=\sum_{\sigma^{(q)}\in S_{q}}a_{q}(\sigma^{(q-1)},\sigma^{(q)})$
$=1.$
Let $a:S_{q-1}\to\mathbb{Q}\setminus\mathbb{Q}^{-}$ be a weight map on $S_{q-1}$, and let $a_{q+1}\circ a_{q}$ be the composition of the fractional weight contribution maps.
Then the weight map extends to $S_{q+1}$ by defining
$a:S_{q+1}\longrightarrow\mathbb{Q}\setminus\mathbb{Q}^{-},$
$\sigma^{(q+1)}\longmapsto a(\sigma^{(q+1)}),$
where for each $\sigma^{(q+1)}\in S_{q+1}$,
$a(\sigma^{(q+1)})=\sum_{\sigma^{(q-1)}\in S_{q-1}}(a_{q+1}\circ a_{q})\bigl(\sigma^{(q-1)},\sigma^{(q+1)}\bigr)\,a(\sigma^{(q-1)}).$
This completes the construction of the extension of the weight map to all simplices of the simplicial complex.
3.6  Comparison of the total weights of 0-simplices and facets
In the preceding sections, we extended the weight map to the simplicial complex, thereby assigning a concentration value to each simplex. The resulting weight distribution over the simplicial complex satisfies several structural properties, which we establish below in the form of propositions and theorem.
First, we compare the total weights of $q$-simplices and (q1)(q-1)-simplices.
Proposition 3.2.
Let $1\leq q\leq dim(\Delta)$. The sum of the weights of all $q$-simplices equals the sum of the weights of all (q1)(q-1)-simplices that are not facets of $\Delta$. That is,
$\sum_{j=1}^{n_{q}}a\bigl(\sigma_{j}^{(q)}\bigr)=\sum_{i=1}^{n_{q-1}}\chi_{i}\,a\bigl(\sigma_{i}^{(q-1)}\bigr),$ (2)
where
[…]
and $n_{q}$ denotes the number of $q$-simplices of $\Delta$.
Proof.
We begin by expanding the definition of the weight of each $q$-simplex in terms of the weights of its (q1)(q-1)-faces. By definition, this gives
$\sum_{j=1}^{n_{q}}a(\sigma_{j}^{(q)})=\sum_{j=1}^{n_{q}}\sum_{i=1}^{n_{q-1}}a_{q}(\sigma_{i}^{(q-1)},\sigma_{j}^{(q)})\,a(\sigma_{i}^{(q-1)}).$
Since all summands are non-negative, we may interchange the order of summation. This yields
$\sum_{j=1}^{n_{q}}a(\sigma_{j}^{(q)})=\sum_{i=1}^{n_{q-1}}\left(\sum_{j=1}^{n_{q}}a_{q}(\sigma_{i}^{(q-1)},\sigma_{j}^{(q)})\right)a(\sigma_{i}^{(q-1)}).$
We now examine the inner sum for a fixed (q1)(q-1)-simplex $\sigma_{i}^{(q-1)}$. If $\sigma_{i}^{(q-1)}$ is not a facet of $\Delta$, then it is contained in at least one $q$-simplex, and hence $k_{\sigma_{i}^{(q-1)}}\neq 0$. By the defining property of the map $a_{q}$, we have
$\sum_{j=1}^{n_{q}}a_{q}(\sigma_{i}^{(q-1)},\sigma_{j}^{(q)})=1,$
and therefore
$\sum_{j=1}^{n_{q}}a_{q}(\sigma_{i}^{(q-1)},\sigma_{j}^{(q)})\,a(\sigma_{i}^{(q-1)})=a(\sigma_{i}^{(q-1)}).$
On the other hand, if $\sigma_{i}^{(q-1)}$ is a facet of $\Delta$, then it is not contained in any $q$-simplex, so $k_{\sigma_{i}^{(q-1)}}=0$. In this case,
$\sum_{j=1}^{n_{q}}a_{q}(\sigma_{i}^{(q-1)},\sigma_{j}^{(q)})=0,$
and consequently
$\sum_{j=1}^{n_{q}}a_{q}(\sigma_{i}^{(q-1)},\sigma_{j}^{(q)})\,a(\sigma_{i}^{(q-1)})=0.$
Combining both cases, we may write
$\sum_{j=1}^{n_{q}}a(\sigma_{j}^{(q)})=\sum_{i=1}^{n_{q-1}}\chi_{i}\,a(\sigma_{i}^{(q-1)}),$
where
[…]
This completes the proof. ∎
Next, we derive a corollary expressing the total weight of $q$-simplices in terms of the total weight of (q1)(q-1)-simplices and the weight of (q1)(q-1)-dimensional facets.
Corollary 3.3.
The sum of the weights of all $q$-simplices together with the sum of the weights of the (q1)(q-1)-dimensional facets is equal to the sum of the weights of the (q1)(q-1)-simplices.
Let $1\leq q\leq dim(\Delta)$. Then
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})+\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)}).$ (3)
Proof.
We begin with equation (2) from Proposition 3.2, which states that
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})=\sum_{i=1}^{n_{q-1}}\chi_{i}\,a(\sigma_{i}^{(q-1)}),$
where $\chi_{i}=1$ if $\sigma_{i}^{(q-1)}$ is not a facet of $\Delta$, and $\chi_{i}=0$ if $\sigma_{i}^{(q-1)}$ is a facet.
We now decompose the sum on the right-hand side into contributions from all (q1)(q-1)-simplices and those which are facets. Writing this explicitly, we obtain
$\sum_{i=1}^{n_{q-1}}\chi_{i}\,a(\sigma_{i}^{(q-1)})=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)})-\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma).$
Substituting this expression into the left-hand side yields
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)})-\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma)$ (4)
Rearranging the terms gives,
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})+\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)}),$
which completes the proof. ∎
The following corollary compares the total weight of $q$-simplices with the total weight of 0-simplices.
Corollary 3.4.
Let $1\leq q\leq dim(\Delta)$. The total weight of the $q$-simplices, together with the weights of all facets of dimension less than $q$, is equal to the total weight of the 0-simplices. That is,
$\sum_{i=1}^{n_{q}}a\bigl(\sigma_{i}^{(q)}\bigr)+\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)<q\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{0}}a(\sigma_{i}^{(0)}).$
Proof.
We apply Corollary 3.3 successively for decreasing dimensions, starting from dimension $q$ and proceeding down to dimension $1$.
For $1\leq q\leq dim(\Delta)$, equation (3) gives
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})+\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)}).$
Applying the equation 4 to the right-hand side for dimension q1q-1, we obtain
$\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)})=\sum_{i=1}^{n_{q-2}}a(\sigma_{i}^{(q-2)})-\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-2\end{subarray}}a(\sigma).$
Substituting this expression into the previous equation and proceeding iteratively over simplices of dimensions $q-1,q-2,\dots,1$, we eventually obtain an expression involving only 0-simplices. After the final step, we are left with
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})+\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)<q\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{0}}a(\sigma_{i}^{(0)}),$
which establishes the result. ∎
Finally, we prove a theorem establishing the equality of total weight between the facets and the 0-simplices.
Theorem 3.5.
The total weight of all facets of $\Delta$ is equal to the total weight of the 0-simplices. That is,
$\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{0}}a(\sigma_{i}^{(0)}).$
Proof.
Let $q=dim(\Delta)$ in Corollary 3.4. Since there are no simplices of dimension greater than $q$, every facet of $\Delta$ has dimension less than or equal to $q$.
Therefore, the sum
$\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)\leq q\end{subarray}}a(\sigma)$
is exactly the sum of the weights of all facets of $\Delta$.
Substituting this into the statement of Corollary 3.4, we obtain
$\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{0}}a(\sigma_{i}^{(0)}),$
which completes the proof. ∎
The weight distribution over a simplicial complex implies that the total weight of all facets is equal to the total weight of the 0-simplices. This reflects the interpretation that weight is initially assigned to vertices and is subsequently distributed across facets as higher-dimensional simplices are formed.
A CM in which every simplex carries a concentration value is called a weighted CM. In what follows, CM will always denote a weighted CM, unless explicitly stated otherwise.
3.7  Inference of higher-order simplices
Within the CM framework, we introduce a boundary-based deterministic rule for admitting higher-order simplices into the complex. For $q\geq 2$, the inclusion of a combinatorially feasible $q$-simplex $\sigma^{(q)}$ is determined solely from the weights of its boundary (q1)(q-1)-simplices.
Let
$E_{q}=\{\sigma^{(q)}\in\mathcal{P}(V)\mid\mathcal{B}(\sigma^{(q)})\subseteq S_{q-1}\}$
denote the set of all combinatorially feasible $q$-simplices whose boundary faces are already present in $S_{q-1}$.
Definition 3.5.
Let $2\leq q\leq dim(\Delta)$. An inclusion rule at dimension $q$ is a map
$\Phi_{q}:E_{q}\to\{0,1\}$
such that the family of $q$-simplices admitted in $\Delta$ is defined by
$S_{q}:=\{\sigma^{(q)}\in E_{q}\mid\Phi_{q}(\sigma^{(q)})=1\}.$
For $\sigma^{(q)}\in E_{q}$, define the aggregated boundary weight
$W_{q}(\sigma^{(q)})=\sum_{\tau\in\mathcal{B}(\sigma^{(q)})}a(\tau).$
At dimension $q$, define the reference weight level of (q1)(q-1)-simplices by
$\bar{a}_{q-1}:=\frac{1}{|S_{q-1}|}\sum_{\tau\in S_{q-1}}a(\tau).$
The dimension-dependent threshold is then defined as
$\theta_{q}:=(q+1)\,\bar{a}_{q-1}=\frac{q+1}{|S_{q-1}|}\sum_{\tau\in S_{q-1}}a(\tau).$
We define a specific weight-based inclusion rule by
$\Phi_{q}(\sigma^{(q)})=1\iff W_{q}(\sigma^{(q)})>\theta_{q}.$
The family of admitted $q$-simplices is therefore
$S_{q}=\{\sigma^{(q)}\in E_{q}\mid W_{q}(\sigma^{(q)})>\theta_{q}\}.$
This defines a concentration-based CM model in which higher-order simplices are added through a deterministic threshold criterion.
In the clique complex construction, the existence of a $q$-simplex is determined purely combinatorially: a simplex is present if and only if all of its (q1)(q-1)-faces are present. Equivalently, the vertices of the simplex span a clique in the $1$-skeleton.
In the CM, this combinatorial feasibility condition is retained but refined by weight information. While the presence of all boundary (q1)(q-1)-simplices guarantees clique compatibility, an additional structural constraint is imposed: a $q$-simplex is admitted only if its aggregated boundary weight exceeds the prescribed threshold $\theta_{q}$.
Consequently, the deterministic threshold criterion filters purely combinatorial closure, selecting only those simplices that exhibit sufficient collective boundary support.
4  Centrality indices in the combinatorial metaplex
Having established the axiomatic framework of the CM, defined by a concentration map on the vertex set and its consistent extension to higher-dimensional simplices, we now turn to the quantitative characterisation of its combinatorial structure. The concentration map fundamentally modifies the induced adjacency structure on the simplicial complex, necessitating the definition of centrality indices adapted to the weighted combinatorial framework.
4.1  Background: centrality in weighted graphs
To motivate the definition of centrality in a CM, we recall the development of centrality indices in graph theory, from unweighted graphs to fully weighted graphs. For the graph-based centrality definitions that follow, we reindex the vertex set as $V=\{v_{1},v_{2},\ldots,v_{n}\}$ so that the indexing of vertices aligns naturally with the entries of the adjacency matrix.
  • (a)
    Unweighted graphs: Let $G=(V,E)$ be a simple graph and $V=\{v_{1},v_{2},\cdots,v_{n}\}$ (consider this vertex set only for graph centralities). Two distinct vertices $v_{i},v_{j}\in V$ are said to be adjacent if and only if the unordered pair $\{v_{i},v_{j}\}$ is an element of the edge set $E$:
    $v_{i}\sim v_{j}\iff\{v_{i},v_{j}\}\in E.$
    The primary centrality index is the degree, defined as the cardinality of the neighbourhood of a vertex (that is, the set containing adjacent vertices). In matrix form, the degree of $v_{i}$ is given by the row-sum of the adjacency matrix [23]:
    $k(v_{i})=\sum_{j=1}^{n}a_{ij},$ (5)
    where $A=[a_{ij}]_{n\times n}$ is the binary adjacency matrix with $a_{ij}=1$ if $\{v_{i},v_{j}\}\in E$ and $a_{ij}=0$ otherwise.
  • (b)
    Edge-weighted graphs: Let $G=(V,E,w)$ be a graph equipped with an edge-weight function $w:E\to\mathbb{R}$. The degree centrality is generalised to the vertex strength, defined as the sum of the weights of edges incident to a vertex [8]:
    $s(v_{i})=\sum_{j=1}^{n}w_{ij},$ (6)
    where $W=[w_{ij}]_{n\times n}$ is the weighted adjacency matrix, with $w_{ij}=w(\{v_{i},v_{j}\})$ if $a_{ij}=1$ and $w_{ij}=0$ otherwise. In this formulation, vertices are treated as unweighted elements.
  • (c)
    Vertex-weighted graphs: To incorporate heterogeneity at the level of vertices, centrality definitions have been extended to the class of vertex-weighted graphs $G=(V,E,W)$, where a weight function $W:V\to\mathbb{R}$ assigns a non-negative real value to each vertex [45]. In this framework, the centrality of a vertex $v_{i}$ depends on the weights of its neighbours:
    $C(v_{i})=\sum_{j=1}^{n}W(v_{j})\,a_{ij},$ (7)
    where $A=[a_{ij}]_{n\times n}$ is the binary adjacency matrix.
  • (d)
    Fully weighted graphs: The most general graph and direct precursor to the CM framework is the fully weighted graph $G=(V,E,W,w)$, in which both vertices and edges are weighted by functions $W:V\to\mathbb{R}$ and $w:E\to\mathbb{R}$. The degree centrality in this case couples vertex weights with edge weights:
    $C_{f}(v_{i})=\sum_{j=1}^{n}W(v_{j})\,w_{ij},$ (8)
    where $A_{w}=[w_{ij}]_{n\times n}$ is the weighted adjacency matrix.
    This formulation ensures consistency with the preceding special cases. Imposing edge homogeneity, that is, $w(e)=1$ for all $e\in E$, recovers the degree centrality of vertex-weighted graphs. Imposing vertex homogeneity, that is $W(v)=1$ for all $v\in V$, produces the formulation for edge-weighted graphs. Simultaneously imposing both conditions reduces the index to the classical degree centrality of unweighted graphs.
In this succeeding section, we extend the fully weighted graph centrality model to the CM by replacing scalar vertex weights with simplicial concentrations and edge weights with facet-induced adjacency strengths.
4.2  Adjacency and degree in the CM
In this section, we introduce a family of centrality maps defined on the set of $q$-simplices. Each centrality map assigns to a simplex a real value determined jointly by its incidence relations and the concentration map, and satisfies structural properties compatible with the CM framework. These maps induce a ranking on simplices and provide a quantitative characterisation of their structural roles within the CM, thereby distinguishing the CM from purely topological simplicial complexes.
We focus on centrality indices derived from the incidence structure (degree) and from the shortest-path structure (closeness and harmonic). In the CM framework, these indices are defined with respect to the weighted adjacency relations induced by the concentration map on $\mathcal{K}_{c}$. We now formalise this construction.
4.2.1 Facet-induced adjacency:
Let $\mathcal{F}(\mathcal{K}_{c})$ denote the set of facets of the CM,
$\mathcal{F}(\mathcal{K}_{c})=\left\{\gamma\in\Delta\;\middle|\;\nexists\tau\in\Delta\text{ such that }\gamma\subsetneq\tau\right\},$
and let the set of $q$-simplices be
$S_{q}(\mathcal{K}_{c})=\{\sigma\in\Delta\mid\dim(\sigma)=q\}.$
Henceforth, $\sigma$ denotes a $q$-simplex unless otherwise stated. We formalise adjacency between simplices in terms of their joint containment in a facet. Two distinct $q$-simplices $\sigma_{i},\sigma_{j}\in S_{q}(\mathcal{K}_{c})$ are said to be adjacent if they are both contained in a common facet. Formally, we define the adjacency relation by
$\sigma_{i}\sim_{F}\sigma_{j}\iff\exists\gamma\in\mathcal{F}(\mathcal{K}_{c})\text{ such that }\sigma_{i}\subsetneq\gamma\text{ and }\sigma_{j}\subsetneq\gamma.$
This definition implies that the adjacency between simplices of dimension $q$ is mediated by a facet containing both [37].
Definition 4.1.
The strength of adjacency between two $q$-simplices $\sigma_{i}$ and $\sigma_{j}$ is defined by
$S(\sigma_{i},\sigma_{j})=\max\left\{a(\gamma)\;\middle|\;\gamma\in\mathcal{F}(\mathcal{K}_{c}),\;\sigma_{i}\cup\sigma_{j}\subseteq\gamma\right\}.$
We set $S(\sigma_{i},\sigma_{j})=0$ if $\sigma_{i}$ and $\sigma_{j}$ are not adjacent or if $i=j$.
The strength of adjacency assigns to each adjacent pair the weight of a mediating facet of maximal concentration, thereby providing a quantitative measure of facet-supported coupling between simplices.
We represent adjacency at dimension $q$ by matrices of size $n_{q}\times n_{q}$. In a simplicial complex, adjacency matrices at level $q$ based on upper adjacency were introduced in [24], here we define facet-induced adjacency.
Definition 4.2.
The simplicial adjacency matrix at level $q$ is denoted
$A^{(q)}=[a_{ij}^{(q)}],$
where
$a_{ij}^{(q)}=\begin{cases}1,&\text{if }i\neq j\text{ and }\sigma_{i}\sim_{F}\sigma_{j},\\ 0,&\text{otherwise}.\end{cases}$
When $q=0$ and we restrict to the $1$-skeleton of the simplicial complex, we recover the adjacency matrix defined for the unweighted graph. Inspired by the weighted simplicial adjacency matrix in [37], we define a facet-induced weighted adjacency matrix as follows.
Definition 4.3.
The weighted adjacency matrix at level $q$ is denoted
$A_{w}^{(q)}=[w_{ij}^{(q)}],$
where
$w_{ij}^{(q)}=\begin{cases}S(\sigma_{i},\sigma_{j}),&\text{if }i\neq j\text{ and }\sigma_{i}\sim_{F}\sigma_{j},\\ 0,&\text{otherwise}.\end{cases}$
When $q=0$ and we restrict to the $1$-skeleton of the simplicial complex, we recover the weighted adjacency matrix of a weighted graph. The matrix $A^{(q)}$ encodes the binary adjacency structure of the simplicial complex, while $A_{w}^{(q)}$ encodes the corresponding weighted adjacency determined by the weight map $a$. Together, these matrices distinguish the unweighted combinatorial structure from its weighted counterpart within the CM framework.
Before introducing degree centrality, we first formalise the definition of centrality at dimension $q$ within the combinatorial metaplex framework.
Definition 4.4.
Let $\mathcal{K}_{c}$ be a CM, and let $S_{q}$ denote the set of all $q$-simplices of $\Delta$. A real-valued function $C$, defined as $C:S_{q}\to\mathbb{R}$, is said to be a centrality index on $\mathcal{K}_{c}$ at level $q$ if it induces at least a semi-order on the elements of $S_{q}$.
That is, if $\sigma_{i},\sigma_{j},\sigma_{k}\in S_{q}$ and $\leq$ is the total order relation on $\mathbb{R}$, then:
  1. (i)
    Reflexivity: $C(\sigma_{i})\leq C(\sigma_{i}),\ \forall\sigma_{i}\in S_{q}$.
  2. (ii)
    Transitivity: if $C(\sigma_{i})\leq C(\sigma_{j})$ and $C(\sigma_{j})\leq C(\sigma_{k})$, then $C(\sigma_{i})\leq C(\sigma_{k})$.
  3. (iii)
    Anti-symmetry: if $C(\sigma_{i})\leq C(\sigma_{j})$ and $C(\sigma_{j})\leq C(\sigma_{i})$, then $C(\sigma_{i})=C(\sigma_{j})$.
  4. (iv)
    Comparability: either $C(\sigma_{i})\leq C(\sigma_{j})$ or $C(\sigma_{j})\leq C(\sigma_{i})$.
We define degree centrality based on facet-induced adjacency, incorporating both the number of adjacent simplices and the strength of their coupling.
(a) Simplicial degree $(k(\sigma_{i}))$
The simplicial degree of a $q$-simplex $\sigma_{i}$ is defined as the cardinality of the set of $q$-simplices adjacent to it. Analytically, it is given by the row sum of the simplicial adjacency matrix:
$k(\sigma_{i})=\sum_{j=1}^{n_{q}}a_{ij}^{(q)}.$
This quantity depends only on the incidence structure of the simplicial complex and is independent of the weight map.
(b) Weighted degree $(D(\sigma_{i}))$
Following the formulation in [37] and the fully weighted graph degree in (8), the weighted degree of a $q$-simplex $\sigma_{i}$ is defined by
$D(\sigma_{i})=\sum_{j=1}^{n_{q}}a(\sigma_{j})\,w_{ij}^{(q)}.$
This quantity measures the aggregate contribution of adjacent simplices, weighted by both their simplex weights and the strengths of the facet-mediated adjacencies.
If $\dim(\Delta)=1$ and $q=0$, then the weighted degree definition reduces to the degree centrality of a vertex in a fully weighted graph, as given in (8). Just as centrality in graphs can be expressed as a linear operator acting on a vertex-weight vector, the CM framework admits a higher-order generalisation in which concentration-induced adjacency operators act on weight functions defined over simplices.
(c) Combined degree centrality $(D^{\alpha}(\sigma_{i}))$
To balance the combinatorial contribution of adjacency cardinality with the contribution of adjacency strength, we introduce a one-parameter family of degree centralities.
Let $\alpha\in[0,1]$. The combined degree centrality of a $q$-simplex $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$ is defined by
$D^{\alpha}(\sigma_{i})=\bigl(k(\sigma_{i})\bigr)^{1-\alpha}\cdot\bigl(D(\sigma_{i})\bigr)^{\alpha}.$
The parameter $\alpha$ regulates the relative contribution of adjacency cardinality and facet-induced coupling strength:
  • (i)
    Topological limit: ($\alpha=0$): $D^{\alpha}(\sigma_{i})=k(\sigma_{i})$.
  • (ii)
    Weighted limit: ($\alpha=1$): $D^{\alpha}(\sigma_{i})=D(\sigma_{i})$.
  • (iii)
    Intermediate regime: ($0<\alpha<1$): $D^{\alpha}(\sigma_{i})$ interpolates between the two.
This formulation defines a continuous family of centrality indices on the CM, allowing simplices to be compared with respect to both their combinatorial adjacency and the concentration structure of the facets that support those adjacencies [50, 51].
4.3  Weighted walks and distances
In the CM, connectivity between simplices is determined not only by facet-induced adjacency, but also by sequences of facets that mediate successive adjacencies. Unlike in graphs, where an edge uniquely determines a transition between two vertices, a pair of simplices in a simplicial complex may be jointly contained in multiple facets. Consequently, a walk between simplices is specified by both the sequence of simplices and the choice of mediating facets at each step. We define the weighted walk following the definition in [37].
Definition 4.5.
Let $\sigma_{\mathrm{start}},\sigma_{\mathrm{end}}\in S_{q}(\mathcal{K}_{c})$. A weighted walk $W$ from $\sigma_{\mathrm{start}}$ to $\sigma_{\mathrm{end}}$ is an alternating sequence of $q$-simplices and facets,
$W:\qquad\sigma_{i_{1}},\;\gamma_{1},\;\sigma_{i_{2}},\;\gamma_{2},\;\dots,\;\gamma_{n},\;\sigma_{i_{n+1}},$
such that:
  1. (i)
    $\sigma_{i_{1}}=\sigma_{\mathrm{start}}$ and $\sigma_{i_{n+1}}=\sigma_{\mathrm{end}}$.
  2. (ii)
    For each $1\leq l\leq n$, the facet $\gamma_{l}\in\mathcal{F}(\mathcal{K}_{c})$ satisfies
    $\sigma_{i_{l}}\subsetneq\gamma_{l}\qquad\text{and}\qquad\sigma_{i_{l+1}}\subsetneq\gamma_{l}.$
  3. (iii)
    For each $1\leq l\leq n$, the mediating facet $\gamma_{l}$ satisfies
    $a(\gamma_{l})=S(\sigma_{i_{l}},\sigma_{i_{l+1}}).$
This definition ensures that each consecutive pair of simplices in the walk is connected by a facet of maximal weight among all admissible mediators. Two simplices are said to be connected if there exists a weighted walk between them.
Let $W$ be a weighted walk as defined above. The length of $W$ is given by
$\mathcal{L}(W)=\sum_{l=1}^{n}\frac{1}{a(\gamma_{l})\,a(\sigma_{i_{l+1}})}.$
This definition reflects the modelling assumption that larger weights of mediating facets and destination simplices indicate stronger combinatorial influence in the formation of higher-order structure. Accordingly, the contribution of each step in the walk is taken to be inversely proportional to the product $a(\gamma_{l})\,a(\sigma_{i_{l+1}})$. Under this choice, steps involving larger weights contribute smaller increments to the total length.
Consequently, walks that pass through highly concentrated facets and simplices have smaller total length, ensuring that shortest paths preferentially follow sequences supported by higher concentration. This definition is therefore consistent with the principle that greater boundary concentration facilitates higher-order interaction.
Definition 4.6.
The shortest path distance between two simplices $\sigma_{i},\sigma_{j}\in S_{q}(\mathcal{K}_{c})$ is defined by
$d(\sigma_{i},\sigma_{j})=\min_{W}\mathcal{L}(W),$
where the minimum is taken over all weighted walks $W$ connecting $\sigma_{i}$ and $\sigma_{j}$.
Because walk length is defined as a sum of strictly positive step contributions, the induced shortest-path function satisfies non-negativity and the triangle inequality. Moreover, $d(\sigma_{i},\sigma_{j})=0$ if and only if $\sigma_{i}=\sigma_{j}$. However, because the step length depends on the destination simplex weight, the resulting distance is not symmetric.
Definition 4.7.
Let $\alpha\geq 0$. The combined shortest path distance between $\sigma_{i},\sigma_{j}\in S_{q}(\mathcal{K}_{c})$ is defined by
$d^{\alpha}(\sigma_{i},\sigma_{j})=\min_{W}\left\{\sum_{l=1}^{n}\left(\frac{1}{w^{(q)}_{i_{l}i_{l+1}}\,a(\sigma_{i_{l+1}})}\right)^{\alpha}\right\},$
where
$w^{(q)}_{i_{l}i_{l+1}}=S(\sigma_{i_{l}},\sigma_{i_{l+1}})=a(\gamma_{l}).$
The parameter $\alpha$ regulates the relative contribution of facet weights and simplex weights in the computation of path length:
  1. (i)
    Weighted limit ($\alpha=1$): $d^{\alpha}$ coincides with the distance induced by the length $\mathcal{L}(W)$.
  2. (ii)
    Intermediate regime ($0<\alpha<1$): $d^{\alpha}$ interpolates between purely combinatorial and weight-dominated distances.
  3. (iii)
    Topological limit ($\alpha=0$): $d^{\alpha}$ reduces to the minimum number of steps in a facet-mediated walk.
4.4  Weighted closeness centrality
We consider the adjacency structure on $S_{q}(\mathcal{K}_{c})$ induced by facet-mediated weighted walks.
Definition 4.8.
A connected component of $S_{q}(\mathcal{K}_{c})$ is a maximal subset $\mathcal{D}\subseteq S_{q}(\mathcal{K}_{c})$ such that for every $\sigma_{i},\sigma_{j}\in\mathcal{D}$ there exists a finite weighted walk joining them.
Closeness centrality quantifies how close a simplex is, on average, to all other simplices within its connected component, with respect to the shortest-path distance induced by weighted walks.
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$ and denote by $\mathcal{D}(\sigma_{i})$ the connected component containing $\sigma_{i}$. The farness of $\sigma_{i}$ is defined as
$F(\sigma_{i})=\sum_{\begin{subarray}{c}\sigma_{j}\in\mathcal{D}(\sigma_{i})\\ \sigma_{j}\neq\sigma_{i}\end{subarray}}d(\sigma_{i},\sigma_{j}).$
Definition 4.9.
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The weighted closeness centrality of $\sigma_{i}$ is defined by
$CC(\sigma_{i})=\frac{1}{\displaystyle\sum_{\begin{subarray}{c}\sigma_{j}\in\mathcal{D}(\sigma_{i})\\ \sigma_{j}\neq\sigma_{i}\end{subarray}}d(\sigma_{i},\sigma_{j})}.$
Thus, simplices with smaller aggregate distance to the remaining simplices in their component receive larger centrality values.
In disconnected settings, the reciprocal formulation above becomes degenerate when distances outside the component are infinite. A standard alternative is the harmonic formulation, which naturally accommodates disconnected structures.
Definition 4.10.
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The harmonic weighted closeness centrality of $\sigma_{i}$ is defined by
$HC(\sigma_{i})=\sum_{\begin{subarray}{c}\sigma_{j}\in S_{q}(\mathcal{K}_{c})\\ \sigma_{j}\neq\sigma_{i}\end{subarray}}\frac{1}{d(\sigma_{i},\sigma_{j})},$
where terms with $d(\sigma_{i},\sigma_{j})=\infty$ are taken to contribute 0.
We now introduce a parameter-dependent version based on the combined weighted distance $d^{\alpha}$.
Definition 4.11.
Let $\alpha\geq 0$ and $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The combined weighted closeness centrality of $\sigma_{i}$ is defined by
$CC^{\alpha}(\sigma_{i})=\frac{1}{\displaystyle\sum_{\begin{subarray}{c}\sigma_{j}\in\mathcal{D}(\sigma_{i})\\ \sigma_{j}\neq\sigma_{i}\end{subarray}}d^{\alpha}(\sigma_{i},\sigma_{j})}.$
Analogously, the harmonic combined centrality is defined by
$HC^{\alpha}(\sigma_{i})=\sum_{\begin{subarray}{c}\sigma_{j}\in S_{q}(\mathcal{K}_{c})\\ \sigma_{j}\neq\sigma_{i}\end{subarray}}\frac{1}{d^{\alpha}(\sigma_{i},\sigma_{j})},$
with infinite distances contributing 0.
The parameter $\alpha$ regulates the relative influence of step count and facet-mediated transition strength in the distance evaluation. When $\alpha=0$, all transitions contribute equally and the metric reduces to the minimal number of facet-mediated steps. When $\alpha=1$, distances are determined entirely by the weights assigned to simplices and their mediating facets along the walk. For $0<\alpha<1$, the resulting centrality interpolates between the step-based regime and the fully weight-driven regime.
We note that betweenness centrality is not defined in the present framework. The classical definition of betweenness relies on symmetric shortest-path distances. In this work, the length of a weighted walk is asymmetric, and consequently the induced distance function does not satisfy symmetry. A consistent extension of classical betweenness centrality is therefore not immediate, and the development of an appropriate asymmetric analogue remains a direction for future investigation.
5  Conclusion
In this work, we introduce the combinatorial metaplex as a mathematical framework for modelling higher-order systems in which interaction structure and intrinsic unit properties are intrinsically coupled. In contrast to classical simplicial-complex-based models, where higher-order simplices arise solely from inclusion relations, the combinatorial metaplex integrates vertex-level intrinsic properties directly into the realisation of higher-dimensional interactions. The resulting structure preserves the combinatorial rigour of simplicial modelling while incorporating discrete intrinsic attributes of the constituent units.
A central contribution of the paper is the formal construction of the concentration map. Beginning with a vertex-level concentration assignment, we provide a constructive extension to all simplices that is compatible with the simplicial incidence structure. In particular, we prove that the total concentration of vertices equals the total concentration of facets, establishing a hierarchical conservation principle. This construction clarifies how vertex-level quantities propagate across dimensions and determine the strength and admissibility of higher-order interactions.
Within this framework, higher-order simplices are realised only when their aggregated boundary concentration exceeds a prescribed threshold. This criterion enables the identification of concentration-supported higher-order interactions during reconstruction, distinguishing them from simplices that would arise solely through combinatorial closure, as in clique-based constructions.
Building on the induced concentration structure, we define facet-mediated adjacency for non-facet simplices and introduce corresponding degree, closeness, and harmonic centrality indices. These measures respect the simplicial organisation while incorporating concentration-driven coupling. The associated one-parameter family of adjacency relations interpolates between purely structural connectivity and concentration-induced interaction.
The combinatorial metaplex therefore provides a unified setting in which interaction topology and intrinsic vertex properties jointly determine higher-order organisation. It may be viewed as a structurally constrained subclass of weighted simplicial complexes characterised by vertex-induced propagation and hierarchical conservation across dimensions. This constraint distinguishes it from arbitrary simplex-level weight assignments and ensures coherence between local intrinsic properties and global higher-order structure.
The framework has been developed in a purely theoretical setting. Extending it to empirical datasets and application domains constitutes a natural direction for future research.